# Statistical inference on partially shape-constrained function-on-scalar linear regression models

### v1

@@ 2.3 Estimation: Shape-constrained regression spline
We now consider fitting the functional regression model by means of spline functions for smooth, flexible, and parsimonious estimation of coefficient functions. Let $\phi_{k}(t)$, $k=1,\ldots,d$ denote spline basis functions defined over $[0,1]$ for a given sequence of knots, then we express $\beta_{j}(t)=\sum_{k=1}^{d}c_{jk}\phi_{k}(t)$, where $c_{jk}$ denotes the basis coefficients. Under the shape constraints on $\beta_{j}(t)$ over $\mathcal{I}_{j}$, we can still employ the regression spline approach based on chosen basis functions, such as $I$-splines [23], $C$-splines [18] or general $B$-splines, with corresponding constraints assigned on a subset of basis coefficients $c_{jk}$. While we focus on partially constrained estimation based on $I$- or $C$- splines in this section, estimation via $B$-splines is also discussed in the Supplementary Material S.6.
[23] first introduced $I$-splines for curve estimation under the monotonicity restriction over the entire domain. As illustrated in the left panel of Figure 2, $I$-spline basis functions are piece-wise quadratic, and at each of the knots indicated by dotted vertical lines, there is exactly one basis function with a non-zero slope. Thus, a monotone increasing (decreasing) spline function can be constructed as a non-negative (non-positive) linear combination of $I$-spline basis functions. For readers interested in generating process of $I$-splines, we refer [23] on how piece-wise quadratic $I$-splines are formed from nonnegative $M$-spline family. Then, if the specific aim is on assigning, for example, monotone increasing restriction over a subset of the domain $\mathcal{I}_{j}$, non-negative coefficients condition should apply to a subset of $c_{jk}$ for which corresponding $I$-spline $\phi_{k}(t)$ display positive slopes within the range of $\mathcal{I}_{j}$.
@@
Figure 2: The piece-wise quadratic $I$-spline basis functions and the piece-wise cubic $C$-spline basis functions under the location of equally spaced knots indicated by the symbol ‘$\times$’ along with the dotted vertical lines.
Next, in terms of convexity constraints, [18] proposed $C$-splines by integrating $I$-splines. As shown in the right panel of Figure 2, $C$-splines form convex and piece-wise cubic functions with non-zero second derivatives at each knot, implying that a linear combination of $C$-splines with non-negative (non-positive) basis coefficients ensures the convexity (concavity) of the function estimator. See Section 2 of [18] for details. Similarly, for the convexity constraints over sub-interval(s) $\mathcal{I}_{j}$, we can impose non-negative restrictions to basis coefficients corresponding to splines showing positive second derivates features within $\mathcal{I}_{j}$.
We now apply the shape-restricted spline function technique to the functional regression model framework to estimate functional coefficients under (2.2). Although methodological or theoretical developments are general to other shape constraints, such as convexity or monotone convexity, we elaborate on the estimation under the monotonicity restriction below and refer to remarks for other restrictions. Let $\phi_{k}^{j}(t)$, $k=1,\ldots,d_{j}$, denote piece-wise quadratic $I$-spline basis functions under a given knot sequence, employed to fit $\beta_{j}(t)$, $j=1,\ldots,p$. Especially for partially constrained coefficients $j\in J$, the knot sequence should include lower and upper boundaries of $\mathcal{I}_{j}$. For the case of multiple disjoint sub-intervals form of $\mathcal{I}_{j}$, all boundaries should be a part of the knot sequence. We then estimate the functional regression coefficients by minimizing the following objective function with respect to non-negative conditions assigning on a subset of basis coefficients $c_{jk}$ depending on the set $J$ and $A_{j}$ as,
$\begin{split}\textrm{Minimize}~{}~{}~{}&\sum_{i=1}^{n}\frac{1}{L_{i}}\sum_{\ell=1}^{L_{i}}\Big{[}Y_{i,\ell}-\sum_{j=1}^{p}{X}_{i,j}\big{\{}\sum_{k=0}^{d_{j}}c_{jk}\phi_{k}^{j}(T_{i,\ell})\big{\}}\Big{]}^{2}\\ \textrm{subject to}~{}~{}~{}&c_{jk}\geq 0~{}~{}~{}\textrm{for}~{}~{}\{(j,k):j\in J~{}\textrm{and}~{}k\in A_{j}\},~{}~{}\end{split}$ (2.17)
where $\phi_{0}^{j}(t)$ indicates the intercept function, i.e., $\phi_{0}^{j}(t)=1$, so that $c_{j0}$ represents the location of monotonicity functions of $\beta_{j}(t)$, the dimensionality of splines $d_{j}$ is determined by the number of knots chosen to approximate $\beta_{j}$, the set $J$ from (2.2) includes indices which coefficients are constrained in $H_{0}$, and the set $A_{j}$ is defined as $\big{\{}k\in\{1,\ldots,d_{j}\}:{\phi_{k}^{j}}{{}^{\prime}}(t)>0,~{}\textrm{for}~{}t\in\mathcal{I}_{j}\big{\}}$.
To obtain shape-restricted $\hat{\beta}_{j}(t)$, we first write our regression problem associated with (2.17) as below. We elaborate for the case of regular evaluation grids, $T_{1},\ldots,T_{L}$, for $i=1,\ldots,n$, and $d_{j}=d$ for representation simplicity, but having different $T_{i,\ell}$ or $d_{j}$ does not affect methodological developments. Let $\textrm{vec}\big{(}[\boldsymbol{c}_{j}]_{j=1}^{p}\big{)}=(\boldsymbol{c}_{1}^{\top},\ldots,\boldsymbol{c}_{p}^{\top})^{\top}$ for $\boldsymbol{c}_{j}=(c_{j0},c_{j1},\ldots,c_{jd})^{\top}$, then
$\mathbf{Y}_{i}=(\mathbf{X}_{i}^{\top}\otimes\Phi)\cdot\textrm{vec}\big{(}[\boldsymbol{c}_{j}]_{j=1}^{p}\big{)}+\boldsymbol{\epsilon}_{i},\quad i=1,\ldots,n,$ (2.18)
under $c_{jk}\geq 0$ for certain $(j,k)$ specified in (2.17), where $\Phi$ denotes a $L\times(d+1)$ basis matrix having its $\ell$-th row consisting of $,\big{(}1,\phi_{1}(T_{\ell}),$$\ldots$,… , $\phi_{d}(T_{\ell})\big{)}$ for $\ell=1,\ldots,L$. Then, the coefficient estimation is associated with finding the projection of $nL$-dimensional $(\mathbf{Y}_{1}^{\top}\ldots,\mathbf{Y}_{n}^{\top})^{\top}$ onto the space spanned by columns of $(\mathbf{X}\otimes\Phi)$ with corresponding constraints on $\boldsymbol{c}_{j}$, where $\mathbf{X}$ denotes the $(n\times p)$ design matrix. By letting $\boldsymbol{\sigma}^{h}\in\mathbb{R}^{nL}$, for $h=1,\ldots,p(d+1)$, denote the set of column vectors on $(\mathbf{X}\otimes\Phi)$ and $\mathcal{L}(\cdot)$ represent the linear space spanned by a given set of vectors, we separate $\boldsymbol{\sigma}^{h}$, $h\in H_{1}$, a set of column vectors associated with basis coefficients with no non-negativity constraint, and define $\mathcal{V}=\mathcal{L}\big{(}\{\boldsymbol{\sigma}^{h};~{}h\in H_{1}\}\big{)}$ indicating the linear space spanned by them. We then further obtain a set of generating vectors, denoted as $\boldsymbol{\delta}^{h}$ for $h\in H_{2}$ satisfying $H_{2}=\{1,\ldots,p(d+1)\}\backslash H_{1}$, that are orthogonal to $\mathcal{V}$. That is, $\boldsymbol{\delta}^{h}=\boldsymbol{\sigma}^{h}-\Pi(\boldsymbol{\sigma}^{h}|\mathcal{V})$, for $h\in H_{2}$, where $\Pi$ indicates a projection operator, such as the Gram-Schmidt process. By extending [18], we can characterize the constrained set as
$\mathcal{C}=\big{\{}\boldsymbol{\mu}\in\mathbb{R}^{nL}:~{}~{}\boldsymbol{\mu}=\sum_{h\in H_{1}}b_{h}\boldsymbol{\sigma}^{h}+\sum_{h\in H_{2}}b_{h}\boldsymbol{\delta}^{h},~{}~{}\textrm{where}~{}b_{h}\geq 0,~{}\textrm{for}~{}h\in H_{2}\}.$ (2.19)
Then the projection of $(\mathbf{Y}_{1}^{\top}\ldots,\mathbf{Y}_{n}^{\top})^{\top}$ to $\mathcal{C}$ can be fulfilled by projecting it onto the set of nonnegative linear combinations of $\boldsymbol{\delta}^{h}$, $h\in H_{2}$, called as constraint cone, and onto the $\mathcal{V}$, separately. We refer to Section 2 of [18] for readers interested in more comprehensive descriptions under a nonparametric regression setting. Owing to the fact that the constraint set $\mathcal{C}$ is a closed convex polyhedral cone in $\mathbb{R}^{nL}$, the projection is unique, and we employ the cone projection algorithm of [19] to find a solution, available as the function qprog or coneA in the R package coneproj. We note that the unconstrained estimates $\tilde{\beta}_{j}(t)$ can be calculated by the least square estimates of $\boldsymbol{c}_{j}$ under (2.18).
@@
Next, we investigate rates of convergence for estimated functional coefficients under the constraints. Let $K_{jn}$ denote the number of knots to fit $\beta_{j}(t)$ growing with $n$, $q$ is the order of the spline, and $\|a\|_{L_{2}}$ be the $L_{2}$ norm of a square-integrable function $a(t)$. Under Conditions $(S1)$$(S6)$, deferred in Supplementary Material S.2, and $\lim_{n\rightarrow\infty}K_{n}\log K_{n}/n=0$ for $K_{n}=\text{max}_{1\leq j\leq p}K_{jn}$, Theorem 2 of [14] and Theorem 6.25 of [26] imply the consistency of unconstrained $\tilde{\boldsymbol{\beta}}(t)=\big{(}\tilde{\beta}_{1}(t),\ldots,\tilde{\beta}_{p}(t)\big{)}^{\top}$ written as, $\|\tilde{\beta}_{j}-\beta_{j}\|^{2}_{L_{2}}=O_{p}(K_{n}n^{-1}+K_{n}^{-2q})$, $j=1,\ldots,p.$ Next, we show the consistency of the constrained estimators by adding the condition (S7) specified in Supplementary Material.
@@
When shape constraints assigned to $\beta_{j}$ is true in $H_{0}$, $K_{nj}=O(n^{1/(2q+1)})$, and conditions $(S1)$$(S7)$ are satisfied, the constrained estimator $\widehat{\boldsymbol{\beta}}(t)=\big{(}\hat{\beta}_{1}(t),\ldots,\hat{\beta}_{p}(t)\big{)}^{\top}$ attains the same rate as the unconstrained estimator.
@@
Remark 3 (Convexity shape constraints).
When the convexity constraint is considered in (2.2) instead of monotonicity, we employ $C$-splines and express the coefficient function in the objective function (2.17) as $\beta_{j}(t)=c_{j0,1}\phi_{0,1}(T_{i,\ell})+c_{j0,2}\phi_{0,2}(T_{i,\ell})+\sum_{k=1}^{d_{j}}c_{jk}\phi_{k}^{j}(T_{i,\ell}),$ where $\phi_{0,1}(t)$ denotes the intercept function, $\phi_{0,2}(t)$ is the identity function, i.e., $\phi_{0,2}(T_{i,\ell})=T_{i,\ell}$, and $\phi_{k}^{j}(t)$ denotes $C$-splines under a given sequence of knots. The first two terms without constraints on $c_{j0,1}$ and $c_{j0,2}$ determine the first-order behaviors of functional coefficients. Then, we assign constraints $c_{jk}\geq 0$, for $(j,k)$, where $j\in J$ and $k\in A_{j}$ with the set $A_{j}$ defined as $\{k\in\{1,\ldots,d_{j}\}:{\phi_{k}^{j}}{{}^{\prime\prime}}(t)>0,~{}\textrm{for}~{}t\in\mathcal{I}_{j}\}$. The non-negativity constraints on basis coefficients associated with $\phi_{k}^{j}(t)$ having positive second derivatives within $\mathcal{I}_{j}$ ensures the convexity of functional estimates over $\mathcal{I}_{j}$. For concavity constraint, we apply the same framework but with non-positivity constraints on a chosen set of $c_{jk}$. In this way, we can consider general $H_{0}$ with various types of shape constraints on $\beta_{j}(t)$ over $\mathcal{I}_{j}$, for $j\in J$.
2.4 Test procedure

### v2

@@ 2.3 Estimation: Shape-constrained regression spline
We now consider fitting the functional regression model by means of spline functions for smooth, flexible, and parsimonious estimation of coefficient functions. Let $\phi_{k}^{j}(t)$, $k=1,\ldots,d_{j}$ denote spline basis functions, e.g., piece-wise polynomial bases, to approximate $\beta_{j}(t)$ under a given sequence of knots. Under (1.1), unconstrained regression estimator can be estimated as $\tilde{\beta}_{j}(t)=\sum_{k=1}^{d_{j}}\tilde{c}_{jk}\phi_{k}^{j}(t)$ with coefficients $\tilde{c}_{jk}$, $j=0,\ldots,p$, $k=0,\dots,d_{j}$ minimizing the sum of squares of residuals,
$\mathcal{L}_{n}\big{(}\{\mathbf{c}_{j}\}_{j=1}^{p}\big{)}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{L_{i}}\sum_{\ell=1}^{L_{i}}\bigg{[}Y_{i,\ell}-\sum_{j=1}^{p}{X}_{i,j}\Big{\{}c_{j0}+\sum_{k=1}^{d_{j}}c_{jk}\phi_{k}^{j}(T_{i,\ell})\Big{\}}\bigg{]}^{2}.\\ \$ (2.15)
Here, $\mathbf{c}_{j}=(c_{j0},c_{j1}\cdots,c_{jd_{j}})^{\top}$, where $c_{j0}$ represents the intercept. The details of the unconstrained estimator are thoroughly discussed in [18] and [17]. Under the shape constraints on $\beta_{j}(t)$ over $\mathcal{I}_{j}$, we can still employ the regression spline approach under (2.15), minimizing the sum of squared residuals, but additionally, along with constraints assigned on a subset of basis coefficients $c_{jk}$ depending on chosen spline basis functions $\phi_{k}^{j}(\cdot)$ and the type of partial constraint.
@@
Figure 3: (Left) The piece-wise quadratic $I$-spline basis functions at the top row and their first derivative functions at the bottom row with the location of equally spaced knots indicated by the symbol ‘$\times$’ along with the dotted vertical lines. (Right) The piece-wise cubic $C$-spline basis functions at the top row and their second derivative functions at the bottom row under the location of equally spaced knots.
Partial monotonicity: To impose partial monotonicity (monotone increasing or decreasing) of $\beta_{j}(s)$ under $H_{0,j}$: $\beta_{j}$ is partially monotone on $\mathcal{I}_{j}$, we use first-order differentiable spline basis functions for $\phi_{k}^{j}(t)$, which are constructed based on a specified knot sequence. Examples include, but are not limited to $B$-splines or I-splines [32, 24] of degree of 2, with a detailed explanation of $I$-splines provided shortly. In particular, a sequence of knot for the $j$th coefficient function estimation should include the lower and upper boundaries of $\mathcal{I}_{j}$, ensuring that the monotonicity condition is applied only within $\mathcal{I}_{j}$. For the case of multiple disjoint sub-intervals form of $\mathcal{I}_{j}$, all boundaries should be a part of the knot sequence. If our particular interest is in partial monotone increasing constraint on $\beta_{j}(t)$, we estimate it by minimizing the sum of squares of residuals $\mathcal{L}_{n}(\{\mathbf{c}_{j}\}_{j=0}^{p})$ in (2.15) with respect to non-negative conditions assigning on the first derivative of $j$th coefficient functions over $t\in\mathcal{I}_{j}$, written as,
 $\textrm{~{}~{}subject to~{}~{}~{}}\Sigma_{k=1}^{d_{j}}c_{jk}{\phi_{k}^{j}}{{}^{\prime}}(t)\geq 0~{}~{}~{}\textrm{for}~{}~{}t\in\mathcal{I}_{j}.$ (2.16)
Here ${\phi_{k}^{j}}{{}^{\prime}}(t)$ denotes the first derivative of chosen basis functions. If the partial monotonicity decreasing condition is imposed, we can switch the non-negativity condition in (2.16) to the non-positivity one.
In this study, we specifically examine the adoption of $I$- spline basis functions, introduced for nonparametric regression function estimation under monotonicity constraints. As illustrated in the top left panel of Figure 3, $I$-spline basis functions are piece-wise quadratic, and at each of the knots indicated by dotted vertical lines, there is exactly one basis function with a positive slope. It is also demonstrated in the bottom left plot, which presents derivatives of $I$-splines having one non-negative value at each knot. Thus, a monotone increasing (decreasing) spline function can be constructed as a non-negative (non-positive) linear combination of $I$-spline basis functions. For readers interested in generating process of $I$-splines, we refer [32] on how piece-wise quadratic $I$-splines are formed from nonnegative $M$-spline family. Indeed, their theoretical properties have been extensively explored under various modeling frameworks [24, 28, 26] along with computational algorithms [25].
Under the choice of $I$-spline basis functions for $\phi_{k}^{j}(t)$ with the given set of knot, a non-negativity condition in (2.16) can be translated to the following non-negativity constraint assigning on a subset of basis coefficients $c_{jk}$, written as,
$\textrm{subject to~{}~{}~{}}c_{jk}\geq 0~{}~{}~{}\textrm{for}~{}~{}k\in S_{j},~{}\textrm{where~{}}S_{j}=\big{\{}k\in\{1,\ldots,d_{j}\}:{\phi_{k}^{j}}{{}^{\prime}}(t)\neq 0,~{}\textrm{for}~{}t\in\mathcal{I}_{j}\big{\}}.$ (2.17)
Here, $S_{j}$ is a set consisting of indices on which $I$-splines basis function has a non-zero slope at a given interval $\mathcal{I}_{j}$. Although $B$-splines are commonly used as basis functions in nonparametric regression, $I$-splines, specifically designed to preserve monotonicity, offer coding advantages by simplifying the generic constraints (2.16) into non-negativity condition on a set of basis coefficients (2.17), Nonetheless, we also provide a specific form of restriction on $c_{jk}$ under $B$-spline basis function approach in Section G of the online Supplementary Material.
Partial convexity. When partial convexity constraint is considered on $\beta_{j}(t)$ under $H_{0,j}$: $\beta_{j}$ is partially convex on $\mathcal{I}_{j}$, we employ second differentiable basis functions for $\phi_{k}^{j}(t)$ under a given knot sequence, which contains lower and upper bounds of $\mathcal{I}_{j}$. Under the general choice of basis functions, such as cubic $B$-spines, the regression coefficient functions can be estimated by minimizing $\mathcal{L}_{n}(\{\mathbf{c}_{j}\}_{j=0}^{p})$ in (2.15) with respect to non-negative conditions assigning on the second derivative of $j$th coefficient functions over $t\in\mathcal{I}_{j}$, written as,
$\textrm{subject to}~{}~{}~{}\Sigma_{k=1}^{d_{j}}c_{jk}{\phi_{k}^{j}}{{}^{\prime\prime}}(t)\geq 0~{}~{}~{}\textrm{for}~{}~{}t\in\mathcal{I}_{j},$ (2.18)
where ${\phi_{k}^{j}}{{}^{\prime\prime}}(t)$ denotes the second derivative of chosen basis functions. Or, under the partial concavity restriction, a non-negativity condition is switched to the non-positivity one.
[24] introduced $C$-spline basis functions to enforce convexity (or concavity) constraints in estimation, analogous to the design of $I$-spline basis functions for modeling monotonic functions. While we refer to Section 2 of [24] for generation and comprehensive reviews of $C$-spline basis functions, they are formed by integrating $I$-splines. As shown in the right panel of Figure 3, $C$-splines form convex and piece-wise cubic functions with one non-negative second derivative value at each knot, implying that a linear combination of $C$-splines with non-negative (non-positive) basis coefficients ensures the convexity (concavity) of the function estimator. Under the choice of $C$-spline basis functions for ${\phi_{k}^{j}}(t)$, we rewrite the sum of squares of residuals as $\mathcal{L}_{n}(\{\mathbf{c}_{j}\}_{j=0}^{p})=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{L_{i}}\sum_{\ell=1}^{L_{i}}[Y_{i,\ell}-\sum_{j=1}^{p}{X}_{i,j}\{c_{j0,1}+c_{j0,2}T_{i,\ell}+\sum_{k=1}^{d_{j}}c_{jk}\phi_{k}^{j}(T_{i,\ell})\}]^{2}$, where $c_{j0,1}+c_{j0,2}T_{i,\ell}$, first two terms in the expression of $\beta_{j}(t)$ determines its first-order behaviors of regression coefficients. Then the generic convexity constraint of (2.18) is rewritten as,
$\textrm{subject to}~{}~{}~{}c_{jk}\geq 0~{}~{}~{}\textrm{for}~{}~{}k\in S_{j}~{}~{}\textrm{where~{}}S_{j}=\big{\{}k\in\{1,\ldots,d_{j}\}:{\phi_{k}^{j}}{{}^{\prime\prime}}(t)\neq 0,~{}\textrm{for}~{}t\in\mathcal{I}_{j}\big{\}}.$ (2.19)
In other words, the non-negativity constraints on basis coefficients associated with $\phi_{k}^{j}(t)$ having non-zero second derivatives within $\mathcal{I}_{j}$ ensure the convexity of functional estimates over $\mathcal{I}_{j}$. If the underlying shape restriction indicates both monotone increasing and convex, we add an extra non-negativity restriction on the coefficient for the first-order term, that is, $c_{j0,2}\geq 0$. For concavity constraint, we apply the same framework but with non-positivity constraints $c_{jk}\leq 0$, for $k\in S_{j}$. Similarly, the additional non-positivity condition on $c_{j0,2}$ imposes partial monotone decreasing and concave function estimation. Consistent with the discussion on partial monotonicity conditions, $B$-splines can always be adopted as basis functions, with their corresponding restrictions on basis coefficients.
Partial positivity. Under the partial positivity constraints on $\beta_{j}(t)$, that is, $H_{0,j}$: $\beta_{j}$ is partially positive on $\mathcal{I}_{j}$, the generic constraint is written as $\Sigma_{k=1}^{d_{j}}c_{jk}{\phi_{k}^{j}}(t)\geq 0$ for $t\in\mathcal{I}_{j}$. Under the choice of quadratic $I$-spline basis functions for ${\phi_{k}^{j}}(t)$, the restriction can be specified as
subject $\textrm{subject to}~{}~{}c_{j0}+c_{j1}+\cdots+c_{jk-1}+\frac{1}{2}c_{jk}\geq 0~{}~{}~{}\textrm{for}~{}~{}k\in S_{j}\textrm{~{}~{}where~{}}S_{j}=\big{\{}k\in\{1,\ldots,d_{j}\}:{\phi_{k}^{j}}(t)\neq 0,~{}\textrm{for}~{}t\in\mathcal{I}_{j}\big{\}}.$
For notational simplicity, we assume a sequence of equispaced knots, including the boundaries of $\mathcal{I}_{k}$. However, in the case of non-equispaced knots, the same approach is applied with a minor modification to the weight on $c_{jk}$ in the constraint, accounting for relative distances between the knots. As shown in the top left panel of Figure 3 illustrating the characteristics of $I$-spline basis functions, ensuring the positivity of $\beta_{j}(t)$ over a sub-interval of the domain requires enforcing positivity on the sum of the basis coefficients corresponding to the $I$-spline basis functions with non-zero values in that sub-interval. Alternatively, $B$-splines or other basis functions can be adopted with properly refined corresponding constraints.
Under $I$- or $C$- splines, each coefficient function $\beta_{j}$ is approximated by a basis function expansion, with the coefficients $c_{j}$ constrained to lie within the constraint set. The quadratic programming problem, expressed in terms of $c_{j}$ with constraints, was transformed into a cone projection problem. Detailed steps are available in [25]. This computation specifically involves the projection onto a closed convex cone, with details on the cone space and the concept of cone projection provided in Section D of the Supplementary Material. Here, the cone projection algorithm determines the elements within this space onto which projection may fall. The R function qprog in the package coneproj fulfills this estimation.
@@
Next, we investigate rates of convergence for estimated functional coefficients under the constraints. Let $K_{jn}$ denote the number of knots to fit $\beta_{j}(t)$ growing with $n$, $q$ is the order of the spline, and $\|a\|_{L_{2}}$ be the $L_{2}$ norm of a square-integrable function $a(t)$. Under Conditions $(S1)$$(S7)$, deferred in Section E of the online Supplementary Material, Theorem 2 of [17] and Theorem 6.25 of [35] imply the consistency of unconstrained $\tilde{\boldsymbol{\beta}}(t)=\big{(}\tilde{\beta}_{1}(t),\ldots,\tilde{\beta}_{p}(t)\big{)}^{\top}$ written as, $\|\tilde{\beta}_{j}-\beta_{j}\|^{2}_{L_{2}}=O_{p}(K_{n}n^{-1}+K_{n}^{-2q})$, $j=1,\ldots,p.$ Next, we show the consistency of the constrained estimators by adding the condition (S8) specified in the online Supplementary Material.
@@
When shape constraints assigned to $\beta_{j}$ is true in $H_{0,j}$ and conditions $(S1)$$(S7)$ are satisfied, the constrained estimator $\widehat{\boldsymbol{\beta}}(t)=\big{(}\hat{\beta}_{1}(t),\ldots,\hat{\beta}_{p}(t)\big{)}^{\top}$ attains the same rate as the unconstrained estimator.
@@
2.4 Testing partial shape constraints
$H_{0,j}:\textrm{$\beta_{j}$ is partially $\mathcal{A}_{j}$-shaped on $\mathcal{I}_{j}$},$ (2.20)
for $j\in J$, where $\mathcal{A}_{j}$ designates qualitative shape constraints to be evaluated by users. For example, if we hypothesize that $\beta_{j}$ is monotone increasing on $\mathcal{I}_{j}$, we read (2.20) as “$H_{0,j}:\textrm{$\beta_{j}$ is monotone increasing on $\mathcal{I}_{j}$}$.” Similarly, if one tests a partial convexity hypothesis, (2.20) can also be written as “$H_{0,j}:\textrm{$\beta_{j}$ is convex on $\mathcal{I}_{j}$}$.” In this paper, monotone and convex hypotheses will mainly be exemplified, yet one can extend our framework to general shape constraints, similarly as covered by many others in the literature [23, 24, 36, 27, 34]. We then propose to reject the null hypothesis $H_{0}$ if the test statistic
is significantly large, where $\widehat{\boldsymbol{\beta}}(t)=\big{(}\hat{\beta}_{1}(t),\ldots,\hat{\beta}_{p}(t)\big{)}^{\top}$ and $\widetilde{\boldsymbol{\beta}}(t)=\big{(}\tilde{\beta}_{1}(t),\ldots,\tilde{\beta}_{p}(t)\big{)}^{\top}$ denote the estimates of coefficient function $\boldsymbol{\beta}(t)=\big{(}\beta_{1}(t),\ldots,\beta_{p}(t)\big{)}^{\top}$ under the null hypothesis $H_{0}$ and its general alternative $H_{1}$, respectively. The test statistic $D_{n}$ is motivated by [30], a goodness-of-fit based test statistic for validating functional constraints on $\beta_{j}(t)$ via linear operator, similar types of the $L^{2}$-based test statistics were frequently employed in the literature for testing the nullity of functional difference for general scope [38, 44, 43]. Indeed, our proposed method corresponds to the global shape constraints if one sets $\mathcal{I}_{j}=[0,1]$. The following Sections 2.2 and 2.3 elaborate on how to obtain $\widehat{\boldsymbol{\beta}}(t)$ under kernel-smoothing and spline-smoothing approaches, respectively. Based on those estimators, we present a bootstrap test procedure in Section 2.4.